1.     Suppose, the market demand and market supply function of rice are Q­D = 30 – 3P and QS= 20 + 2P respectively. Find out – i) Market price and quantity of rice. ii) If due to the increase in the increase in the income of the consumers, the demand increases to QD= 40 - 3P, what will happen to the market price of potato? iii) If the consumers expect that in near future the price of rice will be decreased, how it will affect the current market price of rice? Use graph if required. iv) If the rice farmers of Dinajour send most of their rice to Dhaka City for higher price, what may happen to the market price of rice in Dinajpur? Use graphs. v) If the price becomes Tk. 4, what will happen to the market? vi) If the price becomes Tk. 2, what will happen to the market?



1
Expert's answer
2021-10-13T15:44:39-0400

(i)

"Q_D=30-3P"

"Q_S=20+2P"

"Q_D=Q_S"

"30-3P=20+2P"

"P=2" .

(ii)

"Q_D=40-3P"

"Q_S=20+2P"

"Q_D=Q_S"

"40-3P=20+2P"

"P=4"

Market price increases from 2 to 4.

(iii)

If consumers expect a fall in price of rice in the near future, their current demand will fall and hence the current market price of rice will fall.

(iv)

If rice farmers of Dinajour send most of their rice to Dhaka City for higher price, this will result in a decrease in supply of rice in Dinajour and hence the market price of rice in Dinajour will rise.




As shown in the graph, supply curve will shift leftwards. The initial market price is at "P_E" but due to a decrease in supply, the market price shifts to "P_N" .

(v)

At price=4, less rice will be demanded, supply will be cut down.

(vi)

At price=2, more rice will be demanded, supply will be increased.